Circular No. 8618

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 SD (LINEAR) An apparently asteroidal object discovered by the LINEAR project

(cf. MPEC 2005-S17, MPS 142263; discovery observation below) has been found to show a small condensed coma with a fanlike extension to the northwest on CCD frames taken by R. Naves and M. Campas (0.30-m reflector, Barcelona; Oct. 6.9 UT) and by T. Climent (0.25-m reflector, Alginet, Spain; Oct. 15.9), as reported by M. Kidger, Instituto de Astrofísica de Canarias. Kidger adds that an image taken by E. Reina (Barcelona) on Oct. 8.9 also shows a possible short double tail in p.a. 0^{o} and 300^{o}.

2005 UT  α_{2000}  δ_{2000}  Mag.
Sept.16.30101  0^{h}37^{m}55.^{s}15  +44^{o}22′09′′.5  17.3

The following orbital elements are taken from MPEC 2005-U01:

Epoch = 2005 Aug. 18.0 TT

T = 2005 Nov. 16.4290 TT  ω = 91.^{o}1684
e = 0.497274  Ω = 313.9150 }2000.0
q = 1.575269 AU  i = 21.1390

a = 3.133458 AU n^{o} = 0.1776922 P = 5.55 years

COMET P/2005 S2 (SKIFF)

Improved elliptical orbital elements from MPEC 2005-T76:

Epoch = 2006 July 4.0 TT

T = 2006 June 29.0656 TT  ω = 229.^{o}8241
e = 0.196734  Ω = 161.2691 }2000.0
q = 6.398388 AU  i = 3.1408

a = 7.965465 AU n^{o} = 0.0438417 P = 22.5 years

COMET P/2005 S3 (READ)

Improved orbital elements from MPEC 2005-U07:

T = 2006 Jan. 10.507 TT  ω = 140.^{o}958
e = 0.42059  Ω = 273.428 }2000.0
q = 2.84198 AU  i = 3.481

a = 4.90498 AU n^{o} = 0.090729 P = 10.9 years

2005 October 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green